These incidents occurred in Reading during December of 2004:

----120504: At 0131 hours box 24, Orange & Chestnut streets, AOF 225 S.10th street. Large pile of debris in lot of recently demolished dwelling, with extension to a 2&1/2 story frame row dwelling. Fire placed under control at 0151 hours. Companies assigned: E13,1,5,T1,L1,R1,M-2,C-2 & 8,E3(safety Co.)
----121004: At 0441 hours box 16, 6th & Willow streets, AOF 533 S.6th street. 2&1/2 story occupied MOR dwelling, heavy fire 1st & 2nd floors, with extension to exposure 2 (a similar type attached). 2nd alarm struck at 0459, with the fire placed under control at 0618 hours. Companies assigned: 1st alarm: E5,3,1,T1,L1,R1,M-1,C-4,E13(safety Co.). 2nd alarm: E14,11,Sn1,C-1,2,8,& 600.
----SAME DATE: At 2354 hours a still assignment dispatched for an automobile fire @ 500 block of Thorn street. Engine company arrived to find a fully involved stolen min-van with extension to a large 4&1/2 story warehouse. Box 353, Front & Greenwich streets transmitted at 2359, with the fire placed under control at 0010 hours. Companies assigned: Still alarm: E11 & C-4. Box 353: E3,1,L1,T1,R1,M-2,E9(safety Co.)
----121104: At 1810 hours box 2164, Carpenter Technologies building 108 @ Front & Cathedral streets to assist their fire brigade with a fire in a large melting pot inside a large industrial building. Fire placed under control at 1835 hours. Companies assigned: E11,3,9,1,Sn1,L1,R1,M-1,C-5 & 7.
----121204: At 0705 hours, C-5 on the street from numerous rubbish fires in the N.E. section of the city spotted a loom-up from the Jewish Community Center @ 1700 City Line Ave. The chief arrived to find a 1 story 50x50 pavillion on the grounds of the center fully involved and requested a 2 & 1 response. Companies brought the fire under control at 0728. City FM along with ATF & FBI investigating. Companies assigned: E9,11,Sn1,& C-5.
----121304: At 0302 hours box 39, 5th & Washington streets, AOF 141 N.5th street. 3 story MOR mixed occupancy (apts. above a store) with heavy fire showing 2nd & 3rd floors. 2nd alarm struck at 0314 hours. 3rd alarm struck at 0348 hours for relief, with the fire placed under control at 0410. Two firefighters transported with injuries.Companies assigned: 1st alarm: E3,1,5,L1,T1,R1,M-1,C-5 & 7,E11(safety Co.) 2nd alarm: E13,9,Sn1,M-2,C-1,3,6 & 600. 3rd alarm: E8,12,14,M-3,C-2,4,8, & 9.
----121504: At 0345 hours a still assignment dispatched for a propane tank on fire to the outside rear of the IGA supermarket @ 200 W.Buttonwood street. Companies arrived to find a suspicious fire involving the transfer pipes between 2-500 gallon propane tanks. Box 1718 transmitted for the building at 0353 hours. Fire brought under control with 1 handline & 1 master stream at 0408 hours. Companies assigned: Still alarm: E3,L1,& C-8. Box 1718: E11,1,5,T1,R1,M-1,& C-2.
----SAME DATE: At 1115 hours box 124, 4th & Spruce streets, AOF 325 S.4th street. 3 story occupied MOR dwelling, fire 2nd & 3rd floors. Under control time was 1145 hours. Companies assigned: E3,14,13,T1,L1,M-2,C-1,2,& 9.
----SAME DATE: At 1706 hours box 641, 11th & Spring streets, AOF 1132 Spring street. 3 story EOR occupied multiple dwelling,fire in 1st floor rear apartment. Under control time was 1733. Companies assigned: E9,11,1,L1,R1,M-1,C-1 & 4,E5(safety Co.)
----121604: At 1335 hours box 27, 9th & laurel streets, AOF 961 Cotton street (rear). Fire in a 1 story vacant barber shop attached to a 2&1/2 story EOR dwelling. Fire placed under control at 1347. Companies assigned: E13,3,14,T1,L1,M-3,C-1 & 4,E11(safety Co.)
----SAME DATE: At 1642 hours box 428, 9th & Elm streets, AOF 334 N.9th street. 3 story MOR occupied multiple dwelling,fire 1st floor apartment. Under control time was 1655. Companies assigned: E3,9,13,T1,Sn1,R1,M-3,C-1,4, & 9,E1(safety Co.)
----122104: At 1744 hours box 215, 10th & South streets, AOF 621 S.10th street. 2&1/2 story occupied EOR dwelling, heavy fire in the attic. Under control time was 1807. Companies assigned: E13,8,1,T1,L1,R1,M-2,C-4 & 9,E3(safety Co.)
----122304: At 0706 hours a still assignment dispatched for a malfunctioning oil furnace at 1233 Green street. Companies arrived to find heavy smoke showing from the basement of the 2&1/2 story occupied EOR dwelling. Box 48, 13th & Green streets transmitted at 0710, with the fire placed under control at 0724. Companies assigned: Still alarm: E1,L3,& C-9. Box 48: E9,13,T1,R1,M-3,C-4,E3(safety Co.)
----122704: At 0942 hours box 41, 10th & Penn streets, AOF 43 N.10th street. 3 story occupied MOR multiple dwelling, heavy fire showing 1st floor with reports of occupants trapped. 2nd alarm struck on arrival at 0944. Several occupants self-evacuated via fire escapes and all searches were negative.Fire placed under control at 1011 hours. One firefighter & 2 occupants treated for minor injuries. Companies assigned: 1st alarm: E1,3,8,T1,Sn1,R1,M-3,Amb.6,C-9,E9(safety Co.). 2nd alarm: E14,L3,M-1 & 2,C-1,2,8 & 600.
----123104: At 0918 hours box 231, 16th & Cotton streets, AOF 1554 Cotton street. 2&1/2 story occupied MOR dwelling, fire in the basement. Under control time was 0942. Companies assigned: E13,8,3,T1,R1.M-2,C-5 & 7,E12(safety Co.).

To see pictures of some of the December jobs check out : www.readingpafire.com